Sasparilla County is deciding between two health insurance policies. The county has 16 employees.

HealthSavers would require $129610 today, and would then charge $10916 per employee per year. HealthSavers requires a 3-year contract. Note that the charge per employee will begin at the end of year 1 and the final payment will occur at the end of year 3.

Wellness, Inc. requires a 4-year contract. Wellness, Inc. charges $36172 per employee per year, with the first payment occurring at the end of year 1 and the last payment occurring at the end of year 4. The charge will increase by 16% each subsequent year. If the county can enroll all of its employees in a healthy living course, Wellness, Inc. will provide a rebate of $6157 at the end of year 4. Assume that the county is able to enroll all of its employees in the course.

Assume an interest rate of 8%, compounded annually. Comparing the two policies, what is the equivalent uniform annual worth (EUAW) of the HealthSavers plan. (You do not need to calculate the annual worth of the Wellness, Inc. plan.)

Note: If you are completing this problem again, be sure to calculate the annual worth for the correct plan.
Answer #1

Plan : HealthSavers

Current cost=Co=129610

Annual cost=R=10916*16=$174,656

Rate of interest,i=8%

PW of cost=129610+174656(P/A,8%,3)

PW of cost=129610+174656(P/A,8%,3)=129610+174656*2.577097=$579,715.45

EUAC=PW of cost/(P/A,8%,3)=579715.45/2.577097=$224,949.02